OpenStudy (anonymous):

Help!!!!!! Instructions: A cake is put into an oven to bake. The temperature, H, of the cake (in degrees Farenheit) is a function of how long it has been in the oven for t minute. Thus H=f(t). Question 1: Shetch a graph that represents the relation between t and H. Accurately label the axes. Question 2: Suppose that f'(20)=2. What does this mean in terms of the problem? State your answer in a way that would make sense to someone who know nothing about calculus. Include units. Question 3: Rachel is trying to come up with a formula for this function. She is considering the possible formula f(t)=180(1/2)

OpenStudy (anonymous):

can you please help me

OpenStudy (anonymous):

Based on Rachel's formula calculate f(0). Express your answer correct to two decimal places. Include units with your answer. Question 4:Based on Rachels formula f'(0). Express your answer correct to two decimal places include units. Question 5:Is Rachel's formula a reasonable one for this situation? Explain how you arrived at your conclusion? is the rest of the questions

OpenStudy (anonymous):

@e.mccormick can you please help me

OpenStudy (anonymous):

can you help me

OpenStudy (e.mccormick):

Well, how far did you get?

OpenStudy (anonymous):

I am lost I gotten now where I don't know where to begin

OpenStudy (e.mccormick):

Well, how about collecting what they give you?

OpenStudy (anonymous):

ok so where do we start

OpenStudy (e.mccormick):

They have told you a few things. Not many, but a few. Just pull out those few so you know what means what. For example, if I said position y was a function of x and y=f(x) you would need to write down "Position y=f(x)" So what do you write for your temerature one?

OpenStudy (anonymous):

y=H(t)

OpenStudy (e.mccormick):

They told you it would be "H=f(t)" So the imporant part is: Cake temperature in Farenheit H=f(t)

OpenStudy (anonymous):

ok

OpenStudy (e.mccormick):

You should also note that t = time.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):